OpenStudy (anonymous):

Factor. 60y^2 - 27y -60

OpenStudy (anonymous):

y = 1.25, -0.8

OpenStudy (anonymous):

start be taking out a common factor of 3, to get \[3 (20 y^2-9 y-20)\]

OpenStudy (anonymous):

then you can factor as \[3 (4 y-5) (5 y+4)\]

OpenStudy (anonymous):

60 y^2-27 y-60 = 0 Solve the quadratic equation by completing the square: Divide both sides by 60: y^2-(9 y)/20-1 = 0 Add 1 to both sides: y^2-(9 y)/20 = 1 Add 81/1600 to both sides: y^2-(9 y)/20+81/1600 = 1681/1600 Factor the left hand side: (y-9/40)^2 = 1681/1600 Take the square root of both sides: sqrt(y-9/40) = 41/40 Eliminate the absolute value: y-9/40 = -41/40 or y-9/40 = 41/40 Add 9/40 to both sides: y = -4/5 or y-9/40 = 41/40 Add 9/40 to both sides: y = -4/5 or y = 5/4

OpenStudy (anonymous):

i think the question just said "factor"